OpenStudy (anonymous):

Which logarithmic graph can be used to approximate the value of y in the equation 3y = 4? @jim_thompson5910

OpenStudy (anonymous):

A

jimthompson5910 (jim_thompson5910):

I've seen this question before and it seems oddly worded

OpenStudy (anonymous):

B

jimthompson5910 (jim_thompson5910):

So I'm not sure. You would go from \[\Large 3^y = 4\] to \[\Large y = \log_{3}(4)\]

jimthompson5910 (jim_thompson5910):

as for the graphs, I have no idea why they go this route

OpenStudy (anonymous):

I could not figure it out for crud :P I even went to mathway and wolf

jimthompson5910 (jim_thompson5910):

yeah I couldn't figure it out the first time I saw it. I'm not sure what they're looking for exactly.

jimthompson5910 (jim_thompson5910):

my initial guess was to look for the graph that has a root of \[\Large \log_{3}(4) = \frac{\log(4)}{\log(3)} \approx 1.2618595\]

jimthompson5910 (jim_thompson5910):

but I'm not sure